OpenStudy (anonymous):

could y'all help me? 10m-(7m+4)-5m=16 All I really need are the steps.

OpenStudy (gorv):

just open the bracket

OpenStudy (gorv):

10m-7m-+4-5m=16 when 2 sign come togethr

OpenStudy (gorv):

-+ convert to -

OpenStudy (gorv):

10m-7m-4-5m =16 u try rest ...i m here to help

OpenStudy (anonymous):

so it would become -2m-4=16